Grade 12th passAlgebra

Find the value of the following expressions, when x = –1:
2x2 – x – 2

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that x = -1
Then, substitute the value of x in the question
= (2 × (-1)2)–(-1)–2
= (2 × 1) + 1–2
= 2 + 1–2
= 3–2
= 1

Thanks